OpenStudy (anonymous):

which of the following is equal to cos^2(6x)? I am unsure how to manipulate it. the answer options are: a. sin^2 (3x) /4 b. 1 + cos12x /2 c. 1-cos12x /2 d. 1-2cos6x /4 e. none of the above

OpenStudy (agent0smith):

Use the double angle formulas \[\large \cos^2 \theta = \frac{ 1+\cos 2 \theta }{ 2 }\] here, \[\theta = 6x \].

OpenStudy (anonymous):

i know that i have to use those formulas but the 6 is confusing me